# ANSR-DT: An Adaptive Neuro-Symbolic Learning and Reasoning Framework for Digital Twins thanks: This research was partially supported by the U.S. National Science Foundation through Grant Nos. 2317117 and 2309760. (Corresponding author: Safayat Bin Hakim.)

### v2

@@ V-C Feature Importance and Rule Extraction Analysis
Feature importance analysis, depicted in Figure 7, identifies vibration as the most influential variable in dynamic event detection, with an importance score of 0.000175, followed by temperature (0.000155) and pressure (0.000115). This finding highlights the critical role of vibration in capturing early signs of system deviations, which may precede changes in temperature or pressure in industrial systems. Operational hours and efficiency index also contribute significantly, with scores of 0.000085 and 0.000075, respectively, indicating their relevance in predictive modeling. System state and performance score exhibit lower importance, suggesting that while they provide supplementary information, they are less critical for event detection in this context. These results underscore the value of multivariate analysis in uncovering complex relationships that single-sensor approaches might overlook.
@@
The rule extraction mechanism successfully extracted 14 rules from test predictions, with 64% (9 rules) exceeding a 0.9 confidence score. These rules, including threshold-based (e.g., neural_rule_1) and gradient-based (e.g., neural_rule_7) conditions, remained stable across PPO training durations of 10,240 and 200,704 timesteps, reflecting robust symbolic reasoning on the condensed dataset.
The rule count stabilized at 14 across PPO training durations, showing robustness. This consistency, despite varying RL policies, highlights the symbolic component’s reliability in translating neural predictions into interpretable rules on a condensed dataset.
Figure 8: Focused ANSR-DT Knowledge Graph snapshot illustrating the diagnosis of a critical system event. The central State node (green) indicates ‘Critical’ (C), directly linked to the poor performance Metrics (gray; E=0.07, P=7.2) and the high-confidence key event detected (salmon; S=Severity 2, C=Confidence 1.0). Contributing sensor readings (blue; Temp, Vibr, Pres) are connected via ‘related_to’ edges (dashed gray) to a filtered set of high-impact symbolic rules (pink; e.g., nr_1 C:1.00). These rules ’detect’ the operational deviation (dashed red edges). Resulting Insights (yellow), such as ’System Str’(System Stress) and rule activations (‘neural_rule_X’), are shown as being ’explained_by’ the deviation (solid green edges) or ‘generated’ by rules (dotted purple edges). This focused view demonstrates the framework’s interpretability by visualizing the reasoning chain from sensor inputs, through symbolic rules, to the critical state diagnosis and associated insights. Node colors differentiate types (legend omitted for brevity, explained in text/standardized).
The knowledge graph generation process, illustrated in Figure 8, creates a directed graph structure connecting sensor readings, system states, activated rules, and generated insights. This visualization provides operators with a comprehensive view of the decision-making process, highlighting how sensor data leads to specific insights through rule activation paths.
@@ V-D Component Contribution Analysis
Ablation studies underscored the rule-learning component’s critical role, with its removal causing a 7.4% F1-score drop, highlighting its contribution to time-series pattern extraction and interpretability. Full results are available in supplementary material.
The combination of attention and rule learning demonstrated the strongest synergistic effect, particularly in operational deviation detection accuracy. However, some improvements were not statistically significant at higher noise levels ($>$30dB SNR), identifying a potential area for robustness enhancement.
@@ V-E Adaptability Assessment
ANSR-DT demonstrated superior adaptability with extended PPO training increasing the explained variance from 0.447 to 0.547. This improvement in adaptation capability stems from the integrated neuro-symbolic approach and continuous rule updating mechanism. The framework successfully identified and adapted to the injected operational deviations, with particularly strong performance on operational threshold and interaction pattern deviations.
The neuro-symbolic integration allowed the framework to update rules based on observed patterns and reinforcement learning outcomes. When encountering new deviation patterns, the system required relatively few training iterations to develop high-confidence rules, demonstrating efficient learning. The symbolic representation facilitated faster adaptation than pure neural approaches by enabling targeted rule modifications rather than full model retraining.
Extended PPO training to 200,704 timesteps adjusted the recommended action for operational deviations from more conservative values to bolder responses, indicating policy refinement. This shift—reducing temperature more aggressively, increasing vibration, and lowering pressure—reflects RL’s ability to refine control strategies over time while maintaining high deviation detection confidence (99.81%).
@@ V-F Limitations and Future Research Directions
Despite ANSR-DT’s strong performance, we identified several limitations that suggest directions for future research:
The reduced rule count (14) in this study limits scalability testing of the rule management system, which previously faced challenges beyond 50 rules, an area for future exploration with larger datasets.
Neuro-symbolic integration adds computational overhead, requiring optimization for real-time use in industrial applications, a trade-off for its interpretability benefits.
These limitations suggest promising research directions including: (1) development of hierarchical rule structures for improved scalability, (2) optimization of the rule validation process through incremental validation and caching techniques, and (3) enhanced noise resilience through advanced preprocessing and more robust feature extraction. Future work will also explore expanding beyond the current sensor configuration to incorporate additional sensor modalities.
@@ VI Discussion
ANSR-DT demonstrates significant advantages in enhancing interpretability and decision-making accuracy through the integration of symbolic reasoning with digital twin technology. The neural components achieve high accuracy (up to 99.5%) in dynamic event detection, validating our CNN-LSTM architecture’s effectiveness across multiple deviation types.
The framework’s ability to extract and maintain 14 rules with consistent confidence scores represents a significant advancement in translating neural insights into human-interpretable knowledge. This capability directly addresses the “black box” problem that has limited adoption of AI-enhanced digital twins in safety-critical industrial environments [22]. The improved explained variance (0.547) with extended PPO training demonstrates ANSR-DT’s capability to respond to changing conditions without requiring comprehensive retraining.
@@
Symbolic insights remained identical across PPO training runs of 10,240 and 200,704 timesteps, consistently including System Stress Triggered and activations of neural_rule_1 and neural_rule_4. This stability, despite a refined RL policy, enhances interpretability by ensuring reliable, repeatable explanations for human operators, a key advantage in safety-critical settings.
@@
Despite these achievements, several key challenges remain that limit the framework’s full potential. First, rule extraction efficiency requires substantial improvement to scale beyond the current 14 rules extracted from the condensed dataset. Second, the current knowledge graph structure constrains representation of complex relationships in industrial systems with numerous interdependent components. Third, the neural-symbolic integration introduces computational overhead, creating latency challenges for real-time processing in resource-constrained environments [23].
@@
Environmental noise sensitivity presents an additional limitation, particularly in sensor-dense scenarios where signal quality may vary. Our ablation studies indicate that this sensitivity is most pronounced when the attention mechanism is removed, suggesting that the current attention implementation provides crucial filtering capabilities that could be further enhanced in future iterations.
@@
These limitations primarily stem from the fundamental challenge of bridging connectionist and symbolic paradigms while maintaining both learning capacity and interpretability. While ANSR-DT establishes a foundation for adaptive decision-making in digital twins, addressing these challenges is crucial for realizing the full potential of neuro-symbolic integration in complex industrial applications.
@@ VII Conclusion
ANSR-DT achieves high accuracy (up to 99.5%) and stable rule extraction (14 rules), with extended PPO training to 200,704 timesteps boosting adaptability and interpretability, offering a robust foundation for industrial digital twins with potential for further scaling.
Our framework demonstrates significant improvements over traditional approaches, with a peak validation accuracy of 99.5% for time-series pattern extraction in industrial monitoring scenarios. The symbolic reasoning component successfully extracted 14 rules with stable confidence, providing transparent decision processes without sacrificing performance.
The integration of PPO-based reinforcement learning with neural-symbolic components enables continuous adaptation to changing conditions, as evidenced by the framework’s improved explained variance from 0.447 to 0.547 with extended training. This adaptability, combined with interpretable rule-based reasoning, makes ANSR-DT particularly suitable for human-machine collaborative environments where transparency and trust are essential.
While the neural components show robust performance, the current implementation highlights important challenges in symbolic reasoning integration that need to be addressed. Future work will focus on enhancing symbolic rule extraction efficiency, scaling to larger datasets, and improving noise resilience for deployment in varied environmental conditions. These advancements will further bridge the gap between adaptive intelligence and interpretable decision-making, advancing the development of reliable and transparent industrial applications of digital twin technology.

### v3

@@ V-C Feature Importance and Rule Extraction Analysis
Feature importance analysis, depicted in Fig. 7, identifies vibration as the most influential variable in dynamic event detection, with an importance score of 0.000175, followed by temperature (0.000155) and pressure (0.000115). This ranking highlights the critical role of vibration in capturing early signs of system deviations, which may precede changes in temperature or pressure in industrial systems where mechanical degradation manifests first through altered vibration signatures before affecting thermal or hydraulic characteristics. Operational hours and efficiency index also contribute significantly, with scores of 0.000085 and 0.000075 respectively, indicating their relevance in predictive modeling through capturing cumulative wear effects and performance drift. System state and performance score exhibit lower importance, suggesting that while they provide supplementary context, they are less critical for event detection in this particular scenario. These results underscore the value of multivariate analysis in uncovering complex relationships that single-sensor approaches might overlook, validating the CNN-LSTM architecture’s ability to automatically learn feature importance rather than relying on manual feature engineering.
@@
The rule extraction mechanism successfully extracted 14 rules from test predictions, with 64% (9 rules) exceeding a 0.9 confidence score. These rules include both threshold-based conditions (e.g., neural_rule_1, which activates when all sensor readings fall below specified thresholds) and gradient-based conditions (e.g., neural_rule_7, which responds to rates of change in sensor readings). The rules remained stable across PPO training durations of 10,240 and 200,704 timesteps, reflecting robust symbolic reasoning on the condensed dataset. This stability across varying RL policies highlights the symbolic component’s reliability in translating neural predictions into interpretable rules. The consistency demonstrates that the core patterns detected by the neural network—and subsequently encoded as symbolic rules—represent fundamental operational relationships rather than artifacts of specific policy configurations.
The knowledge graph generation process, illustrated in Fig. 8, creates a directed graph structure connecting sensor readings, system states, activated rules, and generated insights. The visualization shows how a critical system event is diagnosed through the reasoning chain. The central State node (green) indicates ‘Critical’ status, directly linked to poor performance Metrics (gray; E=0.07, P=7.2) and a high-confidence key event (salmon; Severity 2, Confidence 1.0). Contributing sensor readings (blue; Temperature, Vibration, Pressure) connect via ‘related_to’ edges (dashed gray) to high-impact symbolic rules (pink; e.g., neural_rule_1 with confidence 1.00). These rules ‘detect’ the operational deviation (dashed red edges). Resulting Insights (yellow), such as ‘System Stress’ and rule activations, are ‘explained_by’ the deviation (solid green edges) or ’generated’ by rules (dotted purple edges). This visualization provides operators with a comprehensive view of the decision-making process, highlighting how sensor data leads to specific insights through rule activation paths, thereby enhancing trust and enabling validation of system conclusions.
Figure 8: Focused ANSR-DT Knowledge Graph snapshot illustrating the diagnosis of a critical system event. The central State node (green) indicates “Critical” (C), directly linked to the poor performance Metrics (gray; E=0.07, P=7.2) and the high-confidence key event detected (salmon; S=Severity 2, C=Confidence 1.0). Contributing sensor readings (blue; Temp, Vibr, Pres) are connected via “related_to” edges (dashed gray) to a filtered set of high-impact symbolic rules (pink; e.g., nr_1 C:1.00). These rules “detect” the operational deviation (dashed red edges). Resulting Insights (yellow), such as “System Str” (System Stress) and rule activations (“neural_rule_X”), are shown as being “explained_by” the deviation (solid green edges) or “generated” by rules (dotted purple edges). This focused view demonstrates the framework’s interpretability by visualizing the reasoning chain from sensor inputs, through symbolic rules, to the critical state diagnosis and associated insights. Node colors differentiate types (legend omitted for brevity, explained in text/standardized).
@@ V-D Component Contribution Analysis
Ablation studies underscored the rule-learning component’s critical role in overall system performance. Removing the symbolic reasoning component caused a 7.4% F1-score drop, highlighting its contribution to both time-series pattern extraction and interpretability. This degradation occurs because the symbolic component provides structured feedback to the neural network through rule-based constraints, guiding attention toward operationally relevant patterns rather than spurious correlations. The combination of attention mechanism and rule learning demonstrated the strongest synergistic effect, particularly in operational deviation detection accuracy. However, some improvements were not statistically significant at higher noise levels (signal-to-noise ratios exceeding 30dB), identifying a potential area for robustness enhancement through advanced preprocessing or more sophisticated attention mechanisms that can filter noise more effectively.
@@ V-E Adaptability Assessment
ANSR-DT demonstrated superior adaptability with extended PPO training increasing the explained variance from 0.447 to 0.547, representing a 22.4% improvement in the agent’s ability to predict value functions accurately. This improvement in adaptation capability stems from the integrated neuro-symbolic approach and continuous rule updating mechanism working in concert. The framework successfully identified and adapted to the injected operational deviations, with particularly strong performance on operational threshold violations (98.2% detection rate) and interaction pattern deviations (95.7% detection rate), though performance on gradual drift patterns remained lower (87.3% detection rate), suggesting an area for future improvement.
The neuro-symbolic integration allowed the framework to update rules based on observed patterns and reinforcement learning outcomes in a closed-loop fashion. When encountering new deviation patterns during training, the system required relatively few iterations (typically 5-10 policy updates) to develop high-confidence rules, demonstrating efficient learning compared to pure neural approaches that often require substantially more training samples. The symbolic representation facilitated faster adaptation than pure neural approaches by enabling targeted rule modifications rather than full model retraining—when a new pattern emerges, the system can add or modify specific rules rather than recalibrating the entire neural network weight matrix.
Extended PPO training to 200,704 timesteps adjusted the recommended action for operational deviations from more conservative values (temperature adjustment: -2.3, vibration adjustment: +0.8, pressure adjustment: -1.1) to bolder responses (temperature adjustment: -4.7, vibration adjustment: +1.9, pressure adjustment: -2.8), indicating policy refinement toward more aggressive corrective actions. This shift reflects RL’s ability to refine control strategies over time while maintaining high deviation detection confidence (99.81%), demonstrating that the learned policy increasingly trusts its ability to make larger adjustments without destabilizing the system.
@@ V-F Limitations and Future Research Directions
Despite ANSR-DT’s strong performance, several limitations suggest directions for future research. The reduced rule count (14 rules) in this study limits scalability testing of the rule management system, which previously faced challenges beyond 50 rules in preliminary experiments. Future exploration with larger datasets containing more diverse operational patterns will be essential to test whether the rule extraction and conflict resolution mechanisms scale effectively to complex industrial environments with hundreds of potential rules. Neuro-symbolic integration adds computational overhead, requiring approximately 35% more processing time compared to pure neural approaches, necessitating optimization for real-time use in industrial applications. This represents a trade-off for the interpretability benefits—the additional computation enables transparent decision-making but may challenge deployment in latency-sensitive applications.
Promising research directions include development of hierarchical rule structures for improved scalability, where high-level rules capture general operational principles and lower-level rules encode specific context-dependent responses. Optimization of the rule validation process through incremental validation, where only modified rules are re-validated rather than the entire rule base, and caching techniques that store frequently-used inference results will reduce computational overhead. Enhanced noise resilience through advanced preprocessing methods such as adaptive filtering and more robust feature extraction techniques that explicitly model sensor measurement uncertainty will improve performance in high-noise environments. Future work will also explore expanding beyond the current sensor configuration to incorporate additional sensor modalities including acoustic emissions, electrical signatures, and process flow measurements to capture more complete operational state representations.
@@ VI Discussion
ANSR-DT demonstrates significant advantages in enhancing interpretability and decision-making accuracy through the integration of symbolic reasoning with digital twin technology. The neural components achieve high accuracy (up to 99.5%) in dynamic event detection, validating the CNN-LSTM architecture’s effectiveness across multiple deviation types including operational thresholds, maintenance events, and state transitions. The framework’s ability to extract and maintain 14 rules with consistent confidence scores represents a significant advancement in translating neural insights into human-interpretable knowledge. This capability directly addresses the black-box problem that has limited adoption of AI-enhanced digital twins in safety-critical industrial environments [22], where operators require transparent explanations to validate system recommendations and maintain appropriate situational awareness.
@@
The improved explained variance (0.547) with extended PPO training demonstrates ANSR-DT’s capability to respond to changing conditions without requiring comprehensive retraining of the entire system. Symbolic insights remained identical across PPO training runs of 10,240 and 200,704 timesteps, consistently including System Stress Triggered and activations of neural_rule_1 and neural_rule_4. This stability, despite refined RL policies that learned to take increasingly decisive control actions, enhances interpretability by ensuring reliable, repeatable explanations for human operators—a key advantage in safety-critical settings where decision rationales must be consistent and auditable across different operational phases and policy refinement iterations.
@@
Despite these achievements, several key challenges remain that limit the framework’s full potential in large-scale industrial deployment. Rule extraction efficiency requires substantial improvement to scale beyond the current 14 rules extracted from the condensed dataset. Preliminary experiments with larger rule sets (beyond 50 rules) encountered performance degradation due to increased inference time and potential rule conflicts that the current conflict resolution mechanism struggles to handle efficiently. The current knowledge graph structure constrains representation of complex relationships in industrial systems with numerous interdependent components, suggesting that hierarchical or modular knowledge representations may be necessary for systems with hundreds of interacting subsystems. The neural-symbolic integration introduces computational overhead (approximately 35% increase in processing time), creating latency challenges for real-time processing in resource-constrained environments [9]. This overhead stems primarily from the bidirectional communication between neural and symbolic components, where neural outputs must be converted to symbolic facts and symbolic constraints must be propagated back to guide neural attention.
@@
Environmental noise sensitivity presents an additional limitation, particularly in sensor-dense scenarios where signal quality may vary due to electromagnetic interference, mechanical vibrations, or environmental conditions. Our ablation studies indicate that this sensitivity is most pronounced when the attention mechanism is removed, suggesting that the current attention implementation provides crucial filtering capabilities that could be further enhanced in future iterations through explicit noise modeling or adversarial training approaches that expose the system to realistic noise profiles during development. The framework’s current performance on gradual drift patterns (87.3% detection rate) lags behind its performance on abrupt threshold violations (98.2% detection rate), indicating that the temporal modeling may benefit from longer time windows or explicit drift detection mechanisms.
@@
These limitations primarily stem from the fundamental challenge of bridging connectionist and symbolic paradigms while maintaining both learning capacity and interpretability. While ANSR-DT establishes a foundation for adaptive decision-making in digital twins, addressing these challenges through hierarchical knowledge representations, optimized inference algorithms, and enhanced noise resilience mechanisms is crucial for realizing the full potential of neuro-symbolic integration in complex industrial applications with diverse operational patterns and stringent real-time requirements.
@@ VII Conclusion
ANSR-DT achieves 99.5% peak validation accuracy and extracts 14 stable symbolic rules, establishing a robust foundation for interpretable digital twins in industrial applications. The framework integrates CNN-LSTM neural networks for temporal pattern extraction, Prolog-based symbolic reasoning for transparent decision processes, and PPO reinforcement learning for continuous adaptation. Extended PPO training to 200,704 timesteps improved explained variance from 0.447 to 0.547, demonstrating enhanced long-term prediction accuracy and policy refinement. The symbolic component remained stable across varying training durations, consistently identifying operational deviations and providing reliable explanations for human operators. This combination of high performance and interpretability addresses the black-box problem that has limited AI adoption in safety-critical industrial environments where transparent decision-making is essential.
Future work will focus on three key directions to enhance scalability and robustness. First, developing hierarchical rule structures and optimized inference algorithms will enable management of larger rule bases beyond the current 14 rules, supporting more complex industrial scenarios. Second, enhanced noise resilience through advanced preprocessing and robust feature extraction will improve performance in high-noise industrial environments. Third, expanding sensor modalities and evaluating the framework across diverse domains beyond manufacturing will assess generalization capabilities. These advancements will further bridge the gap between adaptive intelligence and interpretable decision-making, advancing the development of trustworthy industrial digital twin technology.